# Badminton Birdie-Like Aerodynamic Alignment of Drifting Dust Grains by Subsonic Gaseous Flows in Protoplanetary Disks

### v1

@@ 5.3 Connection to observations
Several disks show polarization that is elliptical in the polarization orientation with HL Tau as the most prominent example seen from the multiwavelength polarization (Stephens et al., 2017; Lin et al., 2024a) and the high angular resolution polarization (Stephens et al., 2023). If polarization is indeed due to birdie-aligned grains, the $\bm{A}$-field should be in the azimuthal direction with a $180^{\circ}$ degeneracy.
@@ 5.4 Caveats and Future Developments
There are several assumptions made in this exploratory work. First, we assume that angular momentum is predominantly around $\bm{e}_{2}$ (the direction of the gas flow-induced torque) when, generally, we might expect angular momentum around the other two axes as well. In the general case, the rotating motion around $e_{2}$ (and $b_{2}$ considered in this paper) will become precession in 3D. For example, if the grain contains significant angular momentum around $\bm{b}_{3}$, $\Gamma_{2}$ will induce precession around $\bm{e}_{3}$. However, we can expect that the damping torque should inevitably diminish the spin and precession at timescales of order $t_{d}$ (which is rather short in the high-density disk environment where the birdie alignment is most applicable) and reach alignment akin to what is explored in this work. Note that a spheroid will not exhibit drag when spinning around its axis of asymmetry, but it should exist for realistic, irregular grains. Another side effect of the assumption is that grains may artificially exhibit polarization simply from the Gold-projection effect in the unlikely limiting case of $\textbf{{M}}=\textbf{{0}}$ as described in Sec. 5.1.
@@ 6 Conclusions
    We derive a formula for the torque of a smooth body by considering subsonic relative motion and specular reflection of gas on each patch of the surface (Eq. 37). We capture the surface properties of the grains relevant for the grain dynamics by $\bm{K}$, L, and M which characterizes the driving torque from gas pressure, the damping torque as the grain spins, and the torque that arises from the existence of $\bm{A}$, respectively. These quantities allow us to quantify the grain rotational motion without predetermining environmental properties, like the gas density or temperature, and the dynamic state, like the orientation or angular velocity.
@@ Acknowledgements
Z.-Y.D.L. acknowledges support from NASA 80NSSC18K1095, the Jefferson Scholars Foundation, the NRAO ALMA Student Observing Support (SOS) SOSPA8-003, the Achievements Rewards for College Scientists (ARCS) Foundation Washington Chapter, the Virginia Space Grant Consortium (VSGC), and UVA research computing (RIVANNA). ZYL is supported in part by NASA 80NSSC20K0533 and NSF AST-2307199. L.W.L. acknowledges support from NSF AST-1910364 and NSF AST-2307844. MFL acknowledges support from the European Research Executive Agency HORIZON-MSCA-2021-SE-01 Research and Innovation programme under the Marie Skłodowska-Curie grant agreement number 101086388 (LACEGAL). MFL is also grateful for the warmth of the Star Formation group of the Institut de Ciències de l’Espai and the Universidad de Barcelona during his 2024 research stay.
@@ Appendix B Derivation of the drag force on a surface element
In this appendix, we will use a distribution of velocities of the gas for a moving observer and derive the transfer of momentum from the gas onto a patch of surface after integrating over the velocity distribution. The derivation is based on Epstein (1924). We will redefine the symbols utilized in this appendix for clarity and the same symbols in the main text may not correspond to the definition here.
@@
$=-mN\Bigg{[}\begin{pmatrix}-\frac{1}{4h}\\ 0\\ 0\end{pmatrix}+\frac{V}{2\sqrt{\pi h}}\begin{pmatrix}-2\alpha\\ \beta\\ \gamma\end{pmatrix}\Bigg{]}.$ (92)
@@
$d\bm{F}$ $\equiv\frac{[\bm{p}^{(i)}_{\sigma,t}-\bm{p}^{(e)}_{\sigma,t}]dtd\sigma}{dt}$ (93)
@@
We can re-express the equation by connecting it to a few physical quantities. $1/2h$ is simply the isothermal sound speed squared $c_{s}^{2}\equiv kT/m$. Also, $2/\sqrt{\pi h}$ is the average speed $v_{\text{th}}\equiv\sqrt{8kT/(\pi m)}$. The quantity $mN$ is the mass density $\rho_{g}$ as defined in the main text. Furthermore, since $V\alpha$ is the velocity component along the $x$-axis, we can express it as $V\alpha=\bm{V}\cdot\bm{n}$. When expressed in the dot product form and applied to $\bm{n}d\sigma$ for Eq. 93, one can quickly see that $(\bm{V}\cdot\bm{n})\bm{n}$ is simply a vector from $\bm{V}$ that is projected onto the normal direction. We can express this relation through a projection tensor $\textbf{{N}}\equiv\bm{n}\bm{n}$ giving $(\bm{V}\cdot\bm{n})\bm{n}=\textbf{{N}}(\bm{V})$. Implementing these new relations, we have
@@
Appendix C Analytical Solutions to the Integrated Constants
In the main text, Eq. 47 defines an integration form that frequently appears in the equations of motion in Sec. 3. The actual quantities that appeared are $E[\bar{z}^{2}]$, $E[1-\bar{z}^{2}]$, and $E[\bar{z}^{2}-\bar{z}^{4}]$. Note that $E[f(\bar{z})+g(\bar{z})]=E[f(\bar{z})]+E[g(\bar{z})]$ where $f$ and $g$ are some functions of $\bar{z}$. The property means we only need to solve for $E[1]$, $E[\bar{z}^{2}]$, and $E[\bar{z}^{4}]$. Here we provide the analytical solutions.
@@ Appendix C Analytical Solutions to the Integrated Constants
$E[\bar{z}^{2}]=\begin{cases}-\frac{\sqrt{1-e^{2}}}{e^{2}}+\frac{\arcsin{(e)}}{e^{3}}&\text{ , when }\bar{a}<1\\ \frac{2}{3}&\text{ , when }\bar{a}=1\\ \frac{\sqrt{1+\varepsilon^{2}}}{\varepsilon^{2}}-\frac{\operatorname{arcsinh}{(\varepsilon)}}{\varepsilon^{3}}&\text{ , when }\bar{a}>1.\\ \end{cases}$ (99)
@@
$E[\bar{z}^{4}]=\begin{cases}-\frac{(2e^{2}+3)\sqrt{1-e^{2}}}{4e^{4}}+\frac{3\arcsin{(e)}}{4e^{5}}&\text{ , when }\bar{a}<1\\ \frac{2}{5}&\text{ , when }\bar{a}=1\\ \frac{(2\varepsilon^{2}-3)\sqrt{1+\varepsilon^{2}}}{4\varepsilon^{4}}+\frac{3\operatorname{arcsinh}{(\varepsilon)}}{4\varepsilon^{5}}&\text{ , when }\bar{a}>1.\\ \end{cases}$ (100)
With Eq. 98, 99, and 100, we can get $E[1-\bar{z}^{2}]=E[1]-E[\bar{z}^{2}]$ and $E[\bar{z}^{2}-\bar{z}^{4}]=E[\bar{z}^{2}]-E[\bar{z}^{4}]$.

### v2

@@ 5.3 Connection to observations
Several disks show polarization that is elliptical in the polarization orientation with HL Tau as the most prominent example seen from the multiwavelength polarization (Stephens et al., 2017; Lin et al., 2024) and the high angular resolution polarization (Stephens et al., 2023). If polarization is indeed due to birdie-aligned grains, the $\bm{A}$-field should be in the azimuthal direction with a $180^{\circ}$ degeneracy.
@@ 5.4 Caveats and Future Developments
There are several assumptions made in this exploratory work. First, we assume that angular momentum is predominantly around $\bm{\hat{e}}_{2}$ (the direction of the gas flow-induced torque) when, generally, we might expect angular momentum around the other two axes as well. In the general case, the rotating motion around $e_{2}$ (and $b_{2}$ considered in this paper) will become precession in 3D. For example, if the grain contains significant angular momentum around $\bm{\hat{b}}_{3}$, $\Gamma_{2}$ will induce precession around $\bm{\hat{e}}_{3}$. However, we can expect that the damping torque should inevitably diminish the spin and precession at timescales of order $t_{d}$ (which is rather short in the high-density disk environment where the birdie alignment is most applicable) and reach alignment akin to what is explored in this work. Note that a spheroid will not exhibit drag when spinning around its axis of asymmetry, but it should exist for realistic, irregular grains. Another side effect of the assumption is that grains may artificially exhibit polarization simply from the Gold-projection effect in the unlikely limiting case of $\textbf{{M}}=\textbf{{0}}$ as described in Sec. 5.1.
@@ 6 Conclusions
    We derive a formula for the torque of a smooth body in the Epstein regime by considering subsonic relative motion and specular reflection of gas on each patch of the surface (Eq. 37). We capture the surface properties of the grains relevant for the grain dynamics by $\bm{K}$, L, and M which characterizes the driving torque from gas pressure, the damping torque as the grain spins, and the torque that arises from the existence of $\bm{A}$, respectively. These quantities allow us to quantify the grain rotational motion without predetermining environmental properties, like the gas density or temperature, and the dynamic state, like the orientation or angular velocity.
@@ Acknowledgements
We thank the anonymous reviewer for the constructive comments that improved the paper. Z.-Y.D.L. acknowledges support from NASA 80NSSC18K1095, the Jefferson Scholars Foundation, the NRAO ALMA Student Observing Support (SOS) SOSPA8-003, the Achievements Rewards for College Scientists (ARCS) Foundation Washington Chapter, the Virginia Space Grant Consortium (VSGC), and UVA research computing (RIVANNA). ZYL is supported in part by NASA 80NSSC20K0533 and NSF AST-2307199. L.W.L. acknowledges support from NSF AST-1910364 and NSF AST-2307844. MFL acknowledges support from the European Research Executive Agency HORIZON-MSCA-2021-SE-01 Research and Innovation programme under the Marie Skłodowska-Curie grant agreement number 101086388 (LACEGAL). MFL is also grateful for the warmth of the Star Formation group of the Institut de Ciències de l’Espai and the Universidad de Barcelona during his 2024 research stay.
@@ Appendix B Derivation of the drag force on a surface element
In this appendix, we assume rarefied, subsonic gas flow (i.e., Epstein regime) and obtain the drag force on a patch of surface. The validity of the Epstein regime in a disk is discussed in Sec. 4.2. We will use a distribution of velocities of the gas for a moving observer and derive the transfer of momentum from the gas onto a patch of surface after integrating over the velocity distribution. The derivation is based on Epstein (1924). We will redefine the symbols utilized in this appendix for clarity and the same symbols in the main text may not correspond to the definition here.
@@
$=-mN\Bigg{[}\begin{pmatrix}-\frac{1}{4h}\\ 0\\ 0\end{pmatrix}+\frac{V}{2\sqrt{\pi h}}\begin{pmatrix}-2\alpha\\ \beta\\ \gamma\end{pmatrix}\Bigg{]}dtd\sigma.$ (93)
@@
$d\bm{F}$ $\equiv\frac{[\bm{p}^{(i)}_{\sigma,t}-\bm{p}^{(e)}_{\sigma,t}]dtd\sigma}{dt}$
@@
We can re-express the equation by connecting it to a few physical quantities. $1/2h$ is simply the isothermal sound speed squared $c_{s}^{2}\equiv kT/m$. Also, $2/\sqrt{\pi h}$ is the average speed $v_{\text{th}}\equiv\sqrt{8kT/(\pi m)}$. The quantity $mN$ is the mass density $\rho_{g}$ as defined in the main text. Furthermore, since $V\alpha$ is the velocity component along the $x$-axis, we can express it as $V\alpha=\bm{V}\cdot\bm{n}$. When expressed in the dot product form and applied to $\bm{n}d\sigma$ for Eq. B, one can quickly see that $(\bm{V}\cdot\bm{n})\bm{n}$ is simply a vector from $\bm{V}$ that is projected onto the normal direction. We can express this relation through a projection tensor $\textbf{{N}}\equiv\bm{n}\bm{n}$ giving $(\bm{V}\cdot\bm{n})\bm{n}=\textbf{{N}}(\bm{V})$. Implementing these new relations, we have
@@
Appendix C Analytical Solutions to the $E$-Integrals
In the main text, Eq. 47 defines an integration form that frequently appears in the equations of motion in Sec. 3. The actual quantities that appeared are $E[x^{2}]$, $E[1-x^{2}]$, and $E[x^{2}-x^{4}]$. Note that $E[f(x)+g(x)]=E[f(x)]+E[g(x)]$ where $f$ and $g$ are some functions of $x$. The property means we only need to solve for $E[1]$, $E[x^{2}]$, and $E[x^{4}]$. Here we provide the analytical solutions.
@@ Appendix C Analytical Solutions to the $E$-Integrals
$E[x^{2}]=\begin{cases}-\frac{\sqrt{1-e^{2}}}{e^{2}}+\frac{\arcsin{(e)}}{e^{3}}&\text{ , when }\bar{a}<1\\ \frac{2}{3}&\text{ , when }\bar{a}=1\\ \frac{\sqrt{1+\varepsilon^{2}}}{\varepsilon^{2}}-\frac{\operatorname{arcsinh}{(\varepsilon)}}{\varepsilon^{3}}&\text{ , when }\bar{a}>1.\\ \end{cases}$ (99)
@@
$E[x^{4}]=\begin{cases}-\frac{(2e^{2}+3)\sqrt{1-e^{2}}}{4e^{4}}+\frac{3\arcsin{(e)}}{4e^{5}}&\text{ , when }\bar{a}<1\\ \frac{2}{5}&\text{ , when }\bar{a}=1\\ \frac{(2\varepsilon^{2}-3)\sqrt{1+\varepsilon^{2}}}{4\varepsilon^{4}}+\frac{3\operatorname{arcsinh}{(\varepsilon)}}{4\varepsilon^{5}}&\text{ , when }\bar{a}>1.\\ \end{cases}$ (100)
With Eq. 98, 99, and 100, we can get $E[1-x^{2}]=E[1]-E[x^{2}]$ and $E[x^{2}-x^{4}]=E[x^{2}]-E[x^{4}]$.